# Quantum amplification and simulation of strong and ultrastrong coupling of light and matter

### v1

@@ Abstract
Notation
Meaning
QED
quantum electrodynamics
QIP
quantum information processing
RWA
rotating-wave approximation
JC
Jaynes–Cummings
USC
ultrastrong coupling
DSC
deep strong coupling
BEC
Bose-Einstein condensate
VQS
variational quantum simulation
VQA
variational quantum algorithm
SQUID
superconducting quantum interference device
$D(\alpha)$
displacement operator with a complex amplitude $\alpha$
$S\left(\xi\right)$
squeezing operator with a complex parameter $\xi$
$\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$, $\sigma_{\pm}$
Pauli operators
$g$
atom-cavity coupling strength
$J$
inter-cavity coupling strength
$\kappa$
photon loss rate
$\gamma$
atomic spontaneous emission rate
$C$
$C=g^{2}/\kappa\gamma$ single-photon cooperativity
$r$
squeezing degree
$\omega_{\rm cav}$
cavity frequency
$\omega_{q}$
atomic transition frequency
@@ 1.1 The importance of light-matter interactions
The quantum interaction of light and matter is a fundamental area of research in physics that spans across its various fields, encompassing: quantum optics (see, e.g., Refs. Scully and Zubairy (1997); Agarwal (2013); Peřina (1991)), photonics Andrews (2015), cavity quantum electrodynamics (QED) Haroche and Raimond (2006), circuit QED You and Nori (2005); Schoelkopf and Girvin (2008); Clarke and Wilhelm (2008); You and Nori (2011); Gu et al. (2017); Krantz et al. (2019); Kockum and Nori (2019); Kjaergaard et al. (2020); Blais et al. (2021), atom optics Meystre (2001); Gardiner and Zoller (2014); Auzinsh et al. (2014), atom optics Meystre (2001); Gardiner and Zoller (2014); Auzinsh et al. (2014), quantum sensing Degen et al. (2017) and quantum metrology Nawrocki (2015), as well as quantum optical technologies Byrnes and Ilo-Okeke (2021), including quantum cryptography, quantum communications, and optical quantum information processing (QIP) Kok and Lovett (2010). Moreover, light-matter interactions have been actively investigated in condensed matter physics, both at the fundamental level concerning, e.g., cavity quantum materials Schlawin et al. (2022), non-equilibrium condensed matter physics with light Le Hur et al. (2016), light-induced effects in three- and lower-dimensional materials or topological materials Bao et al. (2022); but also at the applied level to understand the principles behind devices like lasers, light-emitting diodes, photodetectors, and solar cells. Moreover, we should also mention the importance of light-matter interactions in astronomy (ranging from analyzing emission and absorption spectra to understanding, e.g, black holes, neutron stars, as well as other stellar and interstellar structures and their evolution), quantum chemistry (especially photochemistry) Levine et al. (2009), and quantum biology (ranging from understanding and simulating photosynthesis to controlling the activity of neurons or other cell types with light using the methods of optogenetics) Lambert et al. (2013).
@@ 1.3 Prototype models for studying light-matter interactions
The most popular description of the interaction between a two-level quantum system (such as a real or artificial atom or a qubit) and a single-mode quantized electromagnetic field (a cavity mode) without the rotating-wave approximation (RWA) — a simplification that disregards non-resonant components in light-matter interaction Hamiltonians — is the quantum Rabi model Rabi (1937). This model simplifies to the standard Jaynes–Cummings (JC) model Jaynes and Cummings (1963); Shore and Knight (1993) under the RWA, i.e., when the counter-rotating terms in the Rabi interaction Hamiltonian are negligible. The multi-atom generalizations of the quantum Rabi and JC models are known as the Dicke Dicke (1954) and Tavis–Cummings Tavis and Cummings (1968) models, respectively. Several interaction-amplification methods exist with the goal of simulating the Rabi (or Dicke) model by using the JC (or Tavis–Cummings) models under the RWA together with classical or quantum drives, as described in more detail below.
The quantum Rabi model serves as a prototype closely linked to various fundamental models and emerging phenomena. These encompass the Hopfield Hopfield (1958) and Jahn–Teller Jahn and Teller (1937); Hines et al. (2004); Larson (2008); Meaney et al. (2010); Dereli et al. (2012) representations, as well as renormalization-group models, including the spin-boson Leggett et al. (1987); Le Hur (2012); Leppäkangas et al. (2018) and Kondo Kondo (1964); Le Hur (2012); Snyman and Florens (2015) descriptions. Thus, simulating the quantum Rabi model enables simulating these or many other models.
The quantum Rabi model and its generalizations have lead to a discovery of a diverse range of novel physical effects (like the creation of photons from the quantum vacuum Ciuti et al. (2005); Johansson et al. (2009, 2010); Wilson et al. (2011); Nation et al. (2012); Johansson et al. (2013)), but they have also brought about notable theoretical complexities. Among these challenges, a prominent one is the breakdown of the RWA. As a result, various aspects of the standard quantum-optical theoretical framework require reconsideration to ensure the precise incorporation of all interaction terms inherent to this regime Kockum et al. (2019); Forn-Díaz et al. (2019); Le Boité (2020). Thus, to correctly describe a quantum Rabi-like system when the counter-rotating terms cannot be neglected, the standard formalisms should be generalized to avoid violating various no-go theorems. For example, as discussed in, e.g., Refs. Salmon et al. (2022); Mercurio et al. (2022) and references therein:
@@
    The conventional master equation used in quantum optics does not accurately capture the way a quantum Rabi-like system interacts with its surrounding environment De Liberato et al. (2009); Settineri et al. (2018); Macrì et al. (2022).
@@
    The expected photon output rate is no longer directly linked to the number of photons within the cavity. This means that usual normal-order correlation functions do not correctly describe the photon emission rate in the quantum Rabi and Dicke models. Thus, it is not possible to observe, e.g., a continuous flow of photons from the ground state of the Rabi model, which would apparently imply a perpetual-motion behavior Stassi et al. (2013); Ciuti and Carusotto (2006); Di Stefano et al. (2018). A direct application of the standard formalism could lead to such unphysical results.
@@
    Additionally, these observations seem to contradict the principle of gauge invariance. Thus, one should be very careful in calculating observables to avoid gauge ambiguities Di Stefano et al. (2019); Gustin et al. (2023); Mercurio et al. (2023); Hughes et al. (2023).
@@ 1.4 Ultrastrong and deep strong coupling regimes
The coupling between light and matter, particularly in the context of quantum systems, is often categorized into four regimes defined by the strength (which can be weak, strong, ultrastrong, and deep strong) of the interaction between the two subsystems Kockum et al. (2019); Forn-Díaz et al. (2019); Le Boité (2020). These regimes have important implications for the behavior and properties of the coupled light-matter systems.
The determination of whether the coupling is strong or weak hinges on the comparison between the value of a coupling strength $g$ and the losses (characterized by some damping rates, say $\kappa_{i}$) in the system. Thus, the weak-coupling and strong-coupling regimes are $g\leq\kappa_{i}$ and $g>\kappa_{i}$, respectively. However, ultrastrong coupling (USC) and deep strong coupling (DSC) should not be misunderstood as just stronger coupling, as their characterization does not involve the losses $\kappa_{i}$, but rather juxtaposes the value of $g$ against the frequency of light (say $\omega_{c}$ of a cavity mode) and the transition frequency of the matter (say $\omega_{a}$ of an atom), the uncoupled constituents of the system. In these regimes, the coupling strength becomes comparable to or even larger than the natural frequencies, which implies that the counter-rotating terms in the quantum Rabi Hamiltonian cannot be neglected. Specifically, the USC and DSC regimes are defined by the conditions $\eta\equiv g/\omega_{c,a}>0.1$ and $\eta>1$ Casanova et al. (2010), respectively. Note that this 10 % threshold value for USC is a matter of convention.
@@
Numerous novel effects inherent in the USC and DSC regimes have been theoretically predicted and their various applications have been proposed, including those summarized in Refs. Kockum et al. (2019); Forn-Díaz et al. (2019). In addition to quantum nonlinear optics, quantum optomechanics, and atom optics, which are described in a greater detail in this review, those proposals encompass also various other fields including: QIP Nataf and Ciuti (2011); Romero et al. (2012); Kyaw et al. (2015); Wang et al. (2016, 2017); Xu et al. (2017); Stassi et al. (2017); Stassi and Nori (2018); Stassi et al. (2020), quantum metrology Ruggenthaler et al. (2018); Garbe et al. (2020); Gietka et al. (2022); Salado-Mejía et al. (2021), quantum plasmonics Tame et al. (2013); Benz et al. (2016); Todisco et al. (2018); Munkhbat et al. (2018), quantum field theory Sabín et al. (2010); Di Stefano et al. (2017, 2019); Savasta et al. (2021), polariton-enhanced superconductivity Sentef et al. (2018); Schlawin et al. (2019), metamaterials Scalari et al. (2012, 2013); Bayer et al. (2017), quantum thermodynamics Seah et al. (2018), and quantum chemistry (especially chemistry QED) Galego et al. (2015); Herrera and Spano (2016); Ebbesen (2016); Bennett et al. (2016); Kowalewski et al. (2016); Flick et al. (2017); Martínez-Martínez et al. (2018); Garcia-Vidal et al. (2021); Genet et al. (2021).
@@
Concerning applications of USC for quantum sensing and quantum metrology, we mention novel high-resolution spectroscopy Ruggenthaler et al. (2018), which takes advantage of reduced linewidths and enhanced signal-to-noise ratios achievable in USC setups. Criticality-enhanced metrology in the USC regime has been predicted for the quantum Rabi Garbe et al. (2020), Dicke Gietka et al. (2022), and Hopfield Salado-Mejía et al. (2021) models. In particular, Ref. Salado-Mejía et al. (2021) predicted an improved precision of a thermometric quantum sensor operating at a quantum phase transition. Recent experimental observation of a superradiant phase transition with emergent cat states in a controlled quantum Rabi model Zheng et al. (2023) shows a feasible way of realizing criticality-enhanced metrology in the USC regime. Mechanical states, which enable quantum-enhanced metrology, can be deterministically generated in USC-regime optomechanics, as shown in Ref. Macrí et al. (2016). The experimental approach developed in Ref. Zheng et al. (2023) can also lead to realizing noise-biased cat qubits for fault-tolerant quantum computation in the USC regime.
@@
It is worth noting that QIP often relies on the coherent exchange or transfer of excitations between light and matter, and this pivotal aspect finds significant relevance in both the strong coupling and USC regimes. However, USC proves notably more efficient at such transfer processes by using virtual photons. The realm of QIP benefits extensively from the capabilities of USC systems. Proposals of applications encompass: QIP protocols with dramatically improved coherence times and quantum-operation fidelity Nataf and Ciuti (2011), ultrafast quantum gate operations Romero et al. (2012); Wang et al. (2017); Xu et al. (2017), long-lasting quantum memories Kyaw et al. (2015); Stassi and Nori (2018), holonomic QIP protocols Wang et al. (2016), quantum error-correction codes Stassi et al. (2017), and scalable quantum processors Stassi et al. (2020). Noteworthy advantages span beyond mere reduction in operation times and improved coherence times and gate fidelity: for example, USC also empowers simpler protocols, where the inherent evolution of an USC system supersedes the need for intricate sequences of quantum gates (see, e.g., Ref. Stassi et al. (2017)). Several of these proposed applications utilize entangled ground states and the underlying parity symmetry.
@@
As mentioned above, it is possible to observe entirely new phenomena in the USC or DSC regimes, e.g., the entangled hybrid light-matter ground state (corresponding to a Schrödinger cat of other Schrödinger cat states) of the quantum Rabi model in the DSC regime, which can be considered a new stable state of matter observed experimentally in Ref. Yoshihara et al. (2017a). The most recent experiment of that research group demonstrated another interesting effect in the DSC regime — an extremely large Lamb shift in a multimode QED system Ao et al. (2023).
@@
Various experimental demonstrations of the USC regime (for reviews see Refs. Kockum et al. (2019); Forn-Díaz et al. (2019) have been reported in different systems including: intersubband polaritons Anappara et al. (2009); Günter et al. (2009); Todorov et al. (2010); Jouy et al. (2011); Geiser et al. (2012); Delteil et al. (2012); Askenazi et al. (2014, 2017), superconducting quantum circuits Niemczyk et al. (2010); Forn-Díaz et al. (2010); Baust et al. (2016); Forn-Díaz et al. (2016); Bosman et al. (2017); Chen et al. (2017); Yoshihara et al. (2017a, b, 2018); Ao et al. (2023), Landau polaritons Muravev et al. (2011); Scalari et al. (2012, 2013); Maissen et al. (2014); Bayer et al. (2017); Paravicini-Bagliani et al. (2019); Keller et al. (2020), organic molecules Schwartz et al. (2011); Kéna-Cohen et al. (2013); Gambino et al. (2014); Gubbin et al. (2014); Mazzeo et al. (2014); Barachati et al. (2018); Todisco et al. (2018); Genco et al. (2018); Eizner et al. (2018); Wang et al. (2022), optomechanical systems George et al. (2016); Benz et al. (2016), hybrid superconducting-optomechanical systems Pirkkalainen et al. (2015a). In several experiments, even the DSC regime has been reached Yoshihara et al. (2017a, 2018); Bayer et al. (2017); Mueller et al. (2020); Ao et al. (2023). To date, the highest normalized coupling constant (close to 2) was experimentally achieved in Ref. Mueller et al. (2020) for plasmon polaritons in three-dimensional nanoparticle metallic crystals.
@@
The USC regime presents the potential for inducing and observing various classes of higher-order processes and nonlinear optical phenomena involving only two-level systems and virtual photons Stassi et al. (2017); Kockum et al. (2017a, b); Cirio et al. (2019); Wang et al. (2024), multiphoton quantum Rabi oscillations Garziano et al. (2015), nonclassical state preparation Ashhab and Nori (2010), parity symmetry breaking and Higgs-like mechanism Garziano et al. (2014), bunched-light emission from individual qubits Garziano et al. (2017), conversion of an atomic superposition state into an entangled photonic state Macrì et al. (2018), as well as simultaneous excitation by a single photon of two or more qubits in a single-mode resonator Garziano et al. (2016); Zhao et al. (2017); Wang et al. (2017); Macrì et al. (2020); Tomonaga et al. (2023) or in different resonators in an array of two or three weakly coupled resonators Garziano et al. (2020). Unfortunately, these and other interesting effects predicted for the quantum Rabi model in the USC regime have not been experimentally demonstrated yet, because of the technological challenges mentioned above. We believe that it will be much easier to induce and observe them by quantum simulations.

### v2

@@ Abstract
Notation Meaning
QED quantum electrodynamics
QIP quantum information processing
RWA rotating-wave approximation
JC Jaynes–Cummings
USC ultrastrong coupling
DSC deep strong coupling
BEC Bose-Einstein condensate
VQS variational quantum simulation
VQA variational quantum algorithm
SQUID superconducting quantum interference device
$D(\alpha)$ displacement operator with a complex amplitude $\alpha$
$S\left(\xi\right)$ squeezing operator with a complex parameter $\xi$
$\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$, $\sigma_{\pm}$ Pauli operators
$g$ atom-cavity coupling strength
$J$ inter-cavity coupling strength
$\kappa$ photon loss rate
$\gamma$ atomic spontaneous emission rate
$C$ $C=g^{2}/\kappa\gamma$ single-photon cooperativity
$r$ squeezing degree
$\omega_{\rm cav}$ cavity frequency
$\omega_{q}$ atomic transition frequency
@@ 1.1 The importance of light-matter interactions
The quantum interaction of light and matter is a fundamental area of research in physics that spans across its various fields, encompassing: quantum optics (see, e.g., Refs. [1, 2, 3]), photonics [4], cavity quantum electrodynamics (QED) [5], circuit QED [6, 7, 8, 9, 10, 11, 12, 13, 14], atom optics [15, 16, 17], quantum sensing [18] and quantum metrology [19], as well as quantum optical technologies [20], including quantum cryptography, quantum communications, and optical quantum information processing (QIP) [21]. Moreover, light-matter interactions have been actively investigated in condensed matter physics, both at the fundamental level concerning, e.g., cavity quantum materials [22], non-equilibrium condensed matter physics with light [23], light-induced effects in three- and lower-dimensional materials or topological materials [24]; but also at the applied level to understand the principles behind devices like lasers, light-emitting diodes, photodetectors, and solar cells. Moreover, we should also mention the importance of light-matter interactions in astronomy (ranging from analyzing emission and absorption spectra to understanding, e.g, black holes, neutron stars, as well as other stellar and interstellar structures and their evolution), quantum chemistry (especially photochemistry) [25], and quantum biology (ranging from understanding and simulating photosynthesis to controlling the activity of neurons or other cell types with light using the methods of optogenetics) [26].
@@ 1.3 Prototype models for studying light-matter interactions
The most popular description of the interaction between a two-level quantum system (such as a real or artificial atom or a qubit) and a single-mode quantized electromagnetic field (a cavity mode) without the rotating-wave approximation (RWA) — a simplification that disregards non-resonant components in light-matter interaction Hamiltonians — is the quantum Rabi model [27]. This model simplifies to the standard Jaynes–Cummings (JC) model [28, 29] under the RWA, i.e., when the counter-rotating terms in the Rabi interaction Hamiltonian are negligible. The multi-atom generalizations of the quantum Rabi and JC models are known as the Dicke [30] and Tavis–Cummings [31] models, respectively. Several interaction-amplification methods exist with the goal of simulating the Rabi (or Dicke) model by using the JC (or Tavis–Cummings) models under the RWA together with classical or quantum drives, as described in more detail below.
The quantum Rabi model serves as a prototype closely linked to various fundamental models and emerging phenomena. These encompass the Hopfield [32] and Jahn–Teller [33, 34, 35, 36, 37] representations, as well as renormalization-group models, including the spin-boson [38, 39, 40] and Kondo [41, 39, 42] descriptions. Thus, simulating the quantum Rabi model enables simulating these or many other models.
The quantum Rabi model and its generalizations have lead to a discovery of a diverse range of novel physical effects (like the creation of photons from the quantum vacuum [43, 44, 45, 46, 47, 48]), but they have also brought about notable theoretical complexities. Among these challenges, a prominent one is the breakdown of the RWA. As a result, various aspects of the standard quantum-optical theoretical framework require reconsideration to ensure the precise incorporation of all interaction terms inherent to this regime [49, 50, 51, 52]. Note that this breakdown is not unique to quantized fields, and has also led to many interesting phenomena with classical fields, including, e.g., Bloch-Siegert shift [53, 54, 55], coherent destruction of tunneling [56, 57, 58], and driving-driven tunneling oscillations [59, 60, 61, 62]. In order to correctly describe a quantum Rabi-like system when the counter-rotating terms cannot be neglected, the standard formalisms should be generalized to avoid violating various no-go theorems. For example, as discussed in, e.g., Refs. [63, 64] and references therein:
@@
    The conventional master equation used in quantum optics does not accurately capture the way a quantum Rabi-like system interacts with its surrounding environment [65, 66, 67, 68].
@@
    The expected photon output rate is no longer directly linked to the number of photons within the cavity. This means that usual normal-order correlation functions do not correctly describe the photon emission rate in the quantum Rabi and Dicke models. Thus, it is not possible to observe, e.g., a continuous flow of photons from the ground state of the Rabi model, which would apparently imply a perpetual-motion behavior [69, 70, 71]. A direct application of the standard formalism could lead to such unphysical results.
@@
    Additionally, these observations seem to contradict the principle of gauge invariance. Thus, one should be very careful in calculating observables to avoid gauge ambiguities [72, 73, 74, 75, 76, 77].
@@ 1.4 Ultrastrong and deep strong coupling regimes
The coupling between light and matter, particularly in the context of quantum systems, is often categorized into four regimes defined by the strength (which can be weak, strong, ultrastrong, and deep strong) of the interaction between the two subsystems [49, 50, 51]. These regimes have important implications for the behavior and properties of the coupled light-matter systems.
The determination of whether the coupling is strong or weak hinges on the comparison between the value of a coupling strength $g$ and the losses (characterized by some damping rates, say $\kappa_{i}$) in the system. Thus, the weak-coupling and strong-coupling regimes are $g\leq\kappa_{i}$ and $g>\kappa_{i}$, respectively. However, ultrastrong coupling (USC) and deep strong coupling (DSC) should not be misunderstood as just stronger coupling, as their characterization does not involve the losses $\kappa_{i}$, but rather juxtaposes the value of $g$ against the frequency of light (say $\omega_{c}$ of a cavity mode) and the transition frequency of the matter (say $\omega_{a}$ of an atom), the uncoupled constituents of the system. In these regimes, the coupling strength becomes comparable to or even larger than the natural frequencies, which implies that the counter-rotating terms in the quantum Rabi Hamiltonian cannot be neglected. Specifically, the USC and DSC regimes are defined by the conditions $\eta\equiv g/\omega_{c,a}>0.1$ and $\eta>1$ [78], respectively. Note that this 10 % threshold value for USC is a matter of convention.
@@
Numerous novel effects inherent in the USC and DSC regimes have been theoretically predicted and their various applications have been proposed, including those summarized in Refs. [49, 50]. In addition to quantum nonlinear optics, quantum optomechanics, and atom optics, which are described in a greater detail in this review, those proposals encompass also various other fields including: QIP [79, 80, 81, 82, 83, 84, 85, 86, 87], quantum metrology [88, 89, 90, 91], quantum plasmonics [92, 93, 94, 95, 96], quantum field theory [97, 98, 74, 99], polariton-enhanced superconductivity [100, 101], metamaterials [102, 103, 104], quantum thermodynamics [105], and quantum chemistry (especially chemistry QED) [106, 107, 108, 109, 110, 111, 112, 113, 114, 115].
@@
Concerning applications of USC for quantum sensing and quantum metrology, we mention novel high-resolution spectroscopy [88], which takes advantage of reduced linewidths and enhanced signal-to-noise ratios achievable in USC setups. Criticality-enhanced metrology in the USC regime has been predicted for the quantum Rabi [89], Dicke [90], and Hopfield [91] models. In particular, Ref. [91] predicted an improved precision of a thermometric quantum sensor operating at a quantum phase transition. Recent experimental observation of a superradiant phase transition with emergent cat states in a controlled quantum Rabi model [116] shows a feasible way of realizing criticality-enhanced metrology in the USC regime. Mechanical states, which enable quantum-enhanced metrology, can be deterministically generated in USC-regime optomechanics, as shown in Ref. [117]. The experimental approach developed in Ref. [116] can also lead to realizing noise-biased cat qubits for fault-tolerant quantum computation in the USC regime.
@@
It is worth noting that QIP often relies on the coherent exchange or transfer of excitations between light and matter, and this pivotal aspect finds significant relevance in both the strong coupling and USC regimes. However, USC proves notably more efficient at such transfer processes by using virtual photons. The realm of QIP benefits extensively from the capabilities of USC systems. Proposals of applications encompass: QIP protocols with dramatically improved coherence times and quantum-operation fidelity [79], ultrafast quantum gate operations [80, 83, 84], long-lasting quantum memories [81, 86], holonomic QIP protocols [82], quantum error-correction codes [85], and scalable quantum processors [87]. Noteworthy advantages span beyond mere reduction in operation times and improved coherence times and gate fidelity: for example, USC also empowers simpler protocols, where the inherent evolution of an USC system supersedes the need for intricate sequences of quantum gates (see, e.g., Ref. [85]). Several of these proposed applications utilize entangled ground states and the underlying parity symmetry.
@@
As mentioned above, it is possible to observe entirely new phenomena in the USC or DSC regimes, e.g., the entangled hybrid light-matter ground state (corresponding to a Schrödinger cat of other Schrödinger cat states) of the quantum Rabi model in the DSC regime, which can be considered a new stable state of matter observed experimentally in Ref. [118]. The most recent experiment of that research group demonstrated another interesting effect in the DSC regime — an extremely large Lamb shift in a multimode QED system [119].
@@
Various experimental demonstrations of the USC regime (for reviews, see Refs. [49, 50]) have been reported in different systems including: intersubband polaritons [120, 121, 122, 123, 124, 125, 126, 127, 128], superconducting quantum circuits [129, 130, 131, 132, 133, 134, 118, 135, 136, 137, 119], Landau polaritons [138, 102, 103, 139, 104, 140, 141, 142], organic molecules [143, 144, 145, 146, 147, 148, 94, 149, 150, 151], optomechanical systems [152, 93, 153, 154], magnons [155, 156, 157], quantum dots [158, 159], and hybrid superconducting-optomechanical systems [160]. In several experiments, even the DSC regime has been reached [118, 136, 104, 161, 119]. To date, the highest normalized coupling constant (close to 2) was experimentally achieved in Ref. [161] for plasmon polaritons in three-dimensional nanoparticle metallic crystals.
@@
The USC regime presents the potential for inducing and observing various classes of higher-order processes and nonlinear optical phenomena involving only two-level systems and virtual photons [85, 162, 163, 164, 165], multiphoton quantum Rabi oscillations [166], nonclassical state preparation [167], parity symmetry breaking and Higgs-like mechanism [168], bunched-light emission from individual qubits [169], conversion of an atomic superposition state into an entangled photonic state [170], as well as simultaneous excitation by a single photon of two or more qubits in a single-mode resonator [171, 172, 83, 173, 174] or in different resonators in an array of two or three weakly coupled resonators [175]. Unfortunately, most of these and other interesting effects predicted for the quantum Rabi model in the USC regime have not been experimentally demonstrated yet, because of the technological challenges mentioned above. We believe that it will be much easier to induce and observe them by quantum simulations.